10.0 g of the strong base, KOH, is dissolved in 20
Answer #1

molarity = mass of solute *1000/gram molar mass of solute * volume of solution in ml

            = 10*1000/56*2000 = 0.089M

KOH is strong base

     KOH -------> K+ + OH-

0.089M                   0.089M

[OH-]   = [KOH]

[OH-]   = 0.089M

POH   = -log[OH-]

          = -log0.089

         = 1.0506

PH    = 14-POH

         = 14-1.0506   = 13

a. 13 >>>> answer